OpenStudy (anonymous):

At what point do the lines represented by the equations 2x + y = 3 and -3 x - y =-7 intersect? I just want to know how to do this. So please show me how to do this.

jimthompson5910 (jim_thompson5910):

Add the equations to get 2x + y = 3 -3 x - y = -7 ------------ -1x + 0y = -4 so -x = -4 x = ???

jimthompson5910 (jim_thompson5910):

what do you get when you solve for x

OpenStudy (anonymous):

Where did you get the 0 from>

jimthompson5910 (jim_thompson5910):

y + (-y) adds to 0y

jimthompson5910 (jim_thompson5910):

so the y terms cancel out

OpenStudy (anonymous):

ok makes more sense now. I just realized that I just forgot all about that negative sign next to the seven. That's why I kept getting wrong answer. Thank you though. You were helpful ^.^

jimthompson5910 (jim_thompson5910):

yw

OpenStudy (anonymous):

Wait

jimthompson5910 (jim_thompson5910):

hmm

jimthompson5910 (jim_thompson5910):

whats up

OpenStudy (anonymous):

Can you help me on this 2x^2 + 10x+12. It is asking which of the following is not a factor. A (x+2) B(x+4) C(x+3) D2 and another one. her Which equation represents a line that is perpendicular to y=-2/3x +5 and passes through the point (2,7)

jimthompson5910 (jim_thompson5910):

2x^2 + 10x+12 2(x^2 + 5x+6) how does it factor further?

OpenStudy (anonymous):

What you mean. I don't really get how to do that problem. :P

jimthompson5910 (jim_thompson5910):

what 2 numbers multiply to 6 AND add to 5

OpenStudy (anonymous):

2?

jimthompson5910 (jim_thompson5910):

2 and what else

OpenStudy (anonymous):

3?

jimthompson5910 (jim_thompson5910):

2 and 3, good

jimthompson5910 (jim_thompson5910):

2+3 = 5 2*3 = 6

jimthompson5910 (jim_thompson5910):

so x^2 + 5x + 6 factors to (x+2)(x+3)

OpenStudy (anonymous):

n then the one that is nota factor is b then (X+4)

jimthompson5910 (jim_thompson5910):

good, x+4 is the only one that doesn't fit

OpenStudy (anonymous):

okie dokie thank u

jimthompson5910 (jim_thompson5910):

np

OpenStudy (anonymous):

Which equation represents a line that is perpendicular to y=-2/3x +5 and passes through the point (2,7) For this one I got y = 3/2x+4

jimthompson5910 (jim_thompson5910):

you got it correct

OpenStudy (anonymous):

How do you know it was correct instantly did you plug in the numbers or something

OpenStudy (anonymous):

SOrry I also a lot of other questions I have to ask for my math homework since I have a test tomorrow :P

OpenStudy (anonymous):

I have to ask u questions for my math test :P

jimthompson5910 (jim_thompson5910):

I plugged x = 2 into y = (3/2)x+4 and got y = 7

jimthompson5910 (jim_thompson5910):

yes I'm using a calculator to help me speed things up

OpenStudy (anonymous):

ok then XD ^.^

OpenStudy (anonymous):

hm I thought you were suppose to plug into the original problem that had the negative next to -2/3

jimthompson5910 (jim_thompson5910):

no, you want to focus on the possible answer

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

I have another one what is the y-intercept of the graph of -2x +3y=-6

OpenStudy (anonymous):

Multiple choices A 2 B -2 C 3 D -3

jimthompson5910 (jim_thompson5910):

plug in x = 0 and solve for y

OpenStudy (anonymous):

ohh I kinda remember this one now

jimthompson5910 (jim_thompson5910):

ok what do you get

OpenStudy (anonymous):

wouldn't you get -2?

jimthompson5910 (jim_thompson5910):

you would

jimthompson5910 (jim_thompson5910):

nice work

OpenStudy (anonymous):

Ok have another one which of the following best represents the inequality y > 5/3 x -1 -

OpenStudy (anonymous):

There was suppose to be a dash line under this >

jimthompson5910 (jim_thompson5910):

they want you to graph this?

OpenStudy (anonymous):

Yes

jimthompson5910 (jim_thompson5910):

ok you'll have to graph y = 5/3x - 1 first

jimthompson5910 (jim_thompson5910):

do you know how to do this?

OpenStudy (anonymous):